OpenStudy (anonymous):

reduce 6x+y+4=0 to the intercept form

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!

imqwerty (imqwerty):

intercept form is like this-\[\frac{ x }{ a }+\frac{ y }{ b }=1\]do it like this-\[6x+y+4=0\]\[6x+y=-4\]\[\frac{ 6x }{ -4}+\frac{ y }{ -4 }=\frac{ -4 }{ -4}\]\[\frac{ x }{ \frac{ -2 }{ 3} }+\frac{ y }{-4 }=1\]done :)